# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ Abstract.
In queueing systems, effective scheduling algorithms are essential for optimizing performance in a wide range of modern applications. While the theory of optimal $M/G/1$ scheduling for mean response time is well established, many modern queueing systems operate with multiple servers. Recently, optimal scheduling for the $M/G/k$ queue has been explored in the heavy traffic limit, but much remains unknown about optimal scheduling in the intermediate regime.
In this paper, we give the first framework for proving nontrivial lower bounds on the mean response time of the $M/G/k$ system under arbitrary scheduling policies. These bounds significantly improve upon previous naive lower bounds, particularly for moderate loads. Key to our approach is a new variable-speed queue, which we call the Increasing Speed Queue, which more accurately captures the work completion behavior of multiserver systems. To analyze the expected work of this Increasing Speed Queue, we develop the DiffeDrift method, a novel manner of employing the drift method/BAR approach, by developing test functions via the solutions to a differential equation.
Key words and phrases:
M/G/k, Drift Method, variable speed queue, scheduling, SRPT, response time, multiserver scheduling, lower bound, queueing, BAR approach
@@ 1. Introduction
In queueing systems, effective scheduling algorithms are essential for optimizing performance in a wide range of modern applications. In the analysis of queueing systems, one of the most important performance metrics is the mean response time, denoted as $\mathbb{E}[T]$. An object of particular interest is the optimal scheduling policy, which minimizes the mean response time among all policies. In the single-server setting, optimality of scheduling policy is well understood. However, many modern queueing systems are multiserver systems, such as server farms and cloud computing services. Thus, a deeper understanding of optimal scheduling in the multiserver setting is required.
@@
In the seminal work of Schrage [29], it is proven that the Shortest Remaining Processing Time (SRPT) minimizes the mean response time of jobs in an $M/G/1$ queue. Schrage and Miller [30] also characterize the mean response time of SRPT exactly. The single-server optimality of SRPT has been used to prove optimality results in the $M/G/k$ under extreme load conditions. When the job sizes are known upon arrival, [9] shows that the multiserver SRPT is optimal in the heavy-traffic regime, as the load $\rho$ approaches the capacity of the system. Likewise, when job sizes are unknown, single-server results have been used to prove multiserver optimality. If the job sizes are unknown but drawn from a known distribution, the optimal scheduling policy in the $M/G/1$ is known to be the Gittins index policy [5]. Correspondingly, [32] shows that the multiserver Gittins policy is optimal in the heavy-traffic regime. At the opposite extreme, in the light traffic limit, it is rare for nontrivial scheduling options to be available, so the choice of scheduling policy is less important.
@@
However, much less is known about optimal $M/G/k$ scheduling under moderate loads. This intermediate regime often reflects the conditions faced by many real-world systems. Recently, the SEK policy introduced in [8] is empirically observed to beat SRPT-$k$ in the intermediate regime. The fact that SRPT-$k$ is not optimal raises an important question,
@@
How much improvement beyond SRPT-$k$ is possible under moderate load?
Specifically, we aim to lower bound the mean response time of $M/G/k$ under arbitrary scheduling policies.
@@ 1.1. Challenges in analyzing $M/G/k$ scheduling
The tagged-job method [16] is a classical tool for analyzing single-server queues under many different scheduling policies, including the broad SOAP class policies [31]. However, the tagged-job method breaks down in the analysis of multiserver scheduling. Unlike the single-server setting, the rate at which a multiserver queue completes work varies, making it intractable to quantify the random amount of work encountered by the tagged job in the system.
Another more flexible tool, the drift method [4], likewise encounters additional challenges when applied to queues with variable work completion rate, and has not been employed in multiserver systems with arbitrary scheduling.
As such, many results focus on the heavy-traffic limit, where the work completion behavior of an $M/G/k$ system is nearly identical to that of a single-server queue, allowing the tagged-job method to be reintroduced. For example, [9] proves that multiserver SRPT is heavy-traffic optimal using a multiserver tagged-job method. Unfortunately, such results are not tight when applied scheduling in the intermediate load regime.
In addition, there have been approximation results for $M/G/k$ systems under specific policies, such as the First-Come First-Serve (FCFS) policy (see, e.g., [15] and [14]). There have also been attempts at an exact solution for multiserver queues using matrix analytic methods, but the state of the art is limited to FCFS scheduling and scheduling policies with no more than two priority classes [34].
@@ 1.2. Our contributions
In this paper, we present the first nontrivial lower bounds on the mean response time of any $M/G/k$ system. Our bounds hold for any arbitrary scheduling policy and hold across all system loads. We introduce a novel single-server variable-speed queue, the Increasing Speed Queue (ISQ-$k$). ISQ-$k$ allows us to bridge the multiserver and single-server queues in this challenging intermediate-load environment. Moreover, we develop a novel DiffeDrift extension to the drift method to analyze variable speed queues, including ISQ-$k$.
We develop our lower bounds by leveraging the existing Work-Integral Number Equality (WINE) formula in a novel way [32]. This equality allows us to establish a lower bound on mean response time by bounding the mean relevant work in the system (see Section 10). The WINE formula applies to all systems: arbitrary scheduling, any number of servers, and any load.
Our key tool for lower bounding mean relevant work is a novel variable-speed single-server queue called the Increasing Speed Queue (ISQ-$k$). We prove that the mean total work of the ISQ-$k$ lower bounds the mean total work of any $M/G/k$ system under arbitrary scheduling policies and at all loads (see Section 6). We build on this result to lower bound the mean relevant work of $M/G/k$ systems under arbitrary scheduling policies.
Our analysis of the ISQ-$k$ system is based on the drift method, which relies on the careful selection of test functions [4]. However, most prior applications of drift methods either use standard test functions that are not customized to the system or employ rescaling techniques to eliminate variable rates [19, 1]. Neither of these methods can accommodate complex variable work completion rates and give a tight analysis of the ISQ-$k$ system.
@@
Figure 1. $M/G/2$ mean response time under SRPT-2, compared against naive lower bounds ($M/G/1$/SRPT and $M/G/\infty$, and against our lower bounds. See Section 10 for the definition of WINE 3 and WINE 4 bounds.
@@ 1.3. Improvement upon prior results
We now show the massive improvement of our lower bound over the prior naive lower bounds. In the prior literature, only two lower bounds on mean response time for the $M/G/k$ queue under arbitrary scheduling policies have appeared. These bounds are the mean service time and the resource-pooled single-server SRPT response time [9]. The resource-pooled system is a system which combines all $k$ servers from the $M/G/k$ into one giant server running at speed 1. Because the resource-pooled $M/G/1$ can simulate any $M/G/k$ scheduling policies, resource-pooled SRPT lower bounds the mean response time of the $M/G/k$ under any scheduling policy. The resource-pooled SRPT lower bound has been used to prove prior heavy-traffic bounds [9, 12].
As an illustrative example, consider a two-server setting with an exponential job size distribution, as shown in Figure 1. The blue line represents the mean response time of the resource pooled SRPT system, and the orange line represents the mean service time bound. The empirical mean response time of $M/G/2$-SRPT is given by the dashed green line – while SRPT-2 is known to not be optimal in this setting [13], it gives a useful sense of where the optimal policy may lie. Both lower bounds are observed to be loose in the intermediate regime.
The red line, labeled WINE 2, represents the combination of the mean service time bound and the resource-pooled SRPT bound via the WINE method. This bound significantly improves upon the naive bounds, particularly in the intermediate regime.
To illustrate the additional benefits of ISQ-$k$, we plot our best lower bound, which is represented by the brown line, labeled WINE 4, in Figure 1. With the help of ISQ-$k$, our lower bound offers a major improvement over WINE 2, on top of WINE 2’s improvement over the prior bounds. Collectively, the region where the optimal response might lie has shrunk by more than half across a wide range of loads, from 0.63 to 0.87, and by substantial margins across a much wider range.
Our results apply to any job size distribution, any number of servers, and any load in the $M/G/k$ system, as well as any scheduling policy. For ease of exposition, we start by proving lower bounds in the 2-server setting in Theorem 4.1 before generalizing our results to the $k$-server setting in Theorem 4.2.
@@ 1.4. Outline of paper
  • Section 2: We discuss prior work on multiserver SRPT, multiserver FCFS and the drift method.
@@
  • Section 4: We present our main results and provide a proof sketch.
@@
  • Section 6: We prove that the ISQ-$k$ system lower bounds the total work in the $M/G/k$ system.
  • Section 7: We prove our main results in the $k=2$ server case, lower bounding the $M/G/2$ under arbitrary scheduling policies.
  • Section 8: We discuss the DiffeDrift method, which we use to derive the ISQ-2 test functions, and how to generalize them to the $k$ server system.
  • Section 9: We prove our main results in the general $k$ case, lower bounding $M/G/k$ mean response time under arbitrary scheduling policies.
  • Section 10: We present a framework for lower bounding the response time of $M/G/k$ based on WINE.
@@ 2. Prior Work
2.1. Shortest-remaining processing time
It is well established that the Shortest-Remaining Processing Time (SRPT-1) policy is optimal in minimizing the mean response time in $M/G/1$ queues [29]. Furthermore, [30] characterizes the mean response time under SRPT in the $M/G/1$ exactly.
In a multiserver setting, the SRPT policy, denoted by SRPT-$k$, has been shown to be heavy-traffic optimal, a result proven in [9]. This work provides an upper bound on the mean response time under SRPT-$k$ that is tight at very high loads.
At the opposite extreme, when there are no arrivals, it is well known that the Shortest Job First (SJF) policy is the optimal nonpreemptive policy. Furthermore, [26] proves that no preemptive policy can outperform the optimal nonpreemptive policy. Since SRPT reduces to SJF when there are no arrivals, SRPT is also optimal in that setting.
However, SRPT-$k$ is not always the optimal policy across all load conditions. Notably, the SEK policy, recently introduced in [8], empirically outperforms the response time of SRPT-$k$ under certain conditions. This raises the question of how much further improvement beyond SRPT-$k$ is possible.
The focus of this paper is on establishing lower bounds for the mean response time of $M/G/k$ systems under arbitrary policies, which has been noted as an open problem [7].
@@ 2.2. Multiserver mean response time analysis
The analysis of multiserver FCFS systems, particularly in the context of $M/G/k$-FCFS queues, is relatively well understood. Early work by [23] and [3] provide bounds on the mean response time based on the first two moments of the job size distribution. [17] conducts an exact analysis of the mean response time in $M/Ph/k$-FCFS systems. Since any arbitrary distribution can be closely approximated by a phase-type distribution, their results offer a useful approximation for mean response time in $M/G/k$-FCFS systems.
More recently, [14] demonstrates how utilizing the full sequence of moments of the job size distribution can yield bounds on mean response time. [24] extends Kingman’s bounds [22], to the multiserver $GI/GI/k$ queue and provide the first simple and explicit bounds for mean waiting time that scale with $1/(1-\rho)$. These results allow for bounds on the mean response time in $M/G/k$-FCFS queues across all load levels. In contrast, our work considers a more general setting by allowing for any arbitrary scheduling policies and focuses on establishing lower bounds.
[32] proves that the Gittins-$k$ policy is nearly optimal in the $M/G/k$ under extremely general conditions. Notably, the Gittins-$k$ policy coincides with SRPT-$k$ when there is perfect information about job sizes, which is the setting we consider in this paper. A key result of their work is the introduction of a new formula for calculating the mean response time, known as the Work Integral Number Equality (WINE). WINE provides a method to translate the exact expected relevant work into the exact mean response time, and is applicable to any queueing system under any scheduling policy. Our approach builds on this formula, using WINE as the basis for developing a framework to establish lower bounds on the mean response time.

### v3

@@ Abstract.
In queueing systems, effective scheduling algorithms are essential for optimizing performance. Optimal scheduling for the M/G/k queue has been explored in the heavy traffic limit, but much remains unknown in the intermediate load regime.
In this paper, we give the first framework for proving nontrivial lower bounds on the mean response time of the M/G/k system under arbitrary scheduling policies. Our bounds tighten previous naive lower bounds by more than 60%, yielding significant improvements particularly for moderate loads. Key to our approach is a new variable-speed queue, which more accurately captures the work completion behavior of multiserver systems. To analyze the expected work of this queue, we develop a novel manner of employing the drift method or the BAR approach, by developing test functions via the solutions to a differential equation.
We validate our results numerically for systems with up to 5 servers and a range of job size distributions.
Department of Industrial Engineering and Management Sciences, Northwestern University, Evanston, United States. ziyuanwang2027@u.northwestern.edu
Department of Industrial Engineering and Management Sciences, Northwestern University, Evanston, United States. izzy.grosof@northwestern.edu
@@ 1. Introduction
In queueing systems, effective scheduling algorithms are essential for optimizing performance in a wide range of modern applications. In the analysis of queueing systems, one of the most important performance metrics is the mean response time, denoted as $\mathbb{E}[T]$. An object of particular interest is the optimal scheduling policy, which minimizes the mean response time among all policies. In the single-server setting, optimal scheduling is well understood.
However, many modern queueing systems are multiserver systems. In particular, many practical systems operate with a small number of parallel servers and preemptible workloads with known and predictable sizes. Examples include personal devices such as cell phones or lower-end laptops with 2–8 cores processing short machine learning tasks, a single GPU partitioned across multiple applications, file transfers over a network with a limited number of concurrent connections, and hospital emergency rooms where triage allows urgent cases to preempt routine consults. Thus, a deeper understanding of optimal scheduling in the multiserver setting with preemption and predictable job sizes is required.
All of the above examples involve preemptible jobs. For instance, training a small machine learning task incurs only minor overhead when pausing and storing model parameters and gradient information. File transfers and downloads can be paused and resumed when needed. Preemption in hospital emergency departments is even more straightforward, as urgent cases clearly take precedence over routine ones.
Moreover, in each of these cases, job duration is at least partially predictable in advance. Machine learning tasks often have relatively predictable durations (Hutter2014), as a function of the number and size of predictors as well as the complexity of the model. GPU workloads consist largely of matrix multiplications with with accurately measurable operations per second (NvidiaGPUPerfGuide). In most file transfer settings, file sizes are known. The triage in emergency departments provides some indication of how long a treatment is likely to take.
@@
The seminal work of Schrage1968Proof proves that the Shortest Remaining Processing Time policy (SRPT) minimizes the mean response time of jobs in an $M/G/1$ queue with preemption and known sizes. Schrage1966 also characterize the mean response time of SRPT exactly. The single-server optimality of SRPT has been used to prove optimality results in the $M/G/k$ under asymptotic load conditions. When the job sizes are known upon arrival, Grosof2018 show that the multiserver SRPT is optimal in the heavy-traffic regime, as the load $\rho$ approaches the capacity of the system. Likewise, when job sizes are unknown, single-server results have been used to prove multiserver optimality. If the job sizes are unknown but drawn from a known distribution, the optimal scheduling policy in the $M/G/1$ is known to be the Gittins index policy (Gittins1979). Correspondingly, Scully2020 show that the multiserver Gittins policy is optimal in the heavy-traffic regime. At the opposite extreme, in the light traffic limit, it is rare for nontrivial scheduling options to be available, so the choice of scheduling policy is less important.
@@
However, much less is known about optimal $M/G/k$ scheduling under moderate loads. This intermediate regime often reflects the conditions faced by many real-world systems. Recently, the SEK policy introduced by Grosof2026 is proven to beat SRPT-$k$ in the moderate load regime. The fact that SRPT-$k$ is not optimal raises an important question,
@@
How much improvement beyond SRPT-$k$ and SEK is possible under moderate load?
Specifically, we aim to prove tighter lower bounds on the mean response time of $M/G/k$ than the existing naive lower bounds systems under arbitrary scheduling policies, see Section 4. We focus on the case where job sizes are known in advance, and discuss how to extend our framework to handle unknown or estimated sizes in Section 11.
@@ 1.1. Challenges in analyzing $M/G/k$ scheduling
The tagged-job method (Harchol-Balter2013) is a classical tool for analyzing single-server queues under many different scheduling policies, including the broad SOAP class of policies (Scully2018). However, the tagged-job method breaks down in the analysis of multiserver scheduling. Unlike the single-server setting, the rate at which a multiserver queue completes work varies, making it intractable to quantify the random amount of work encountered by the tagged job in the system.
As such, many results focus on the heavy-traffic limit, where the work completion behavior of an $M/G/k$ system is nearly identical to that of a single-server queue, allowing the tagged-job method to be reintroduced. For example, Grosof2018 prove that multiserver SRPT is heavy-traffic optimal using a multiserver tagged-job method. Unfortunately, such results are not tight when applied to scheduling in the intermediate load regime.
Another more flexible tool, the drift method (Eryilmaz2012), likewise encounters additional challenges when applied to queues with variable work completion rate, and has not been employed in multiserver systems with arbitrary scheduling.
In addition, there have been approximation results for $M/G/k$ systems under specific policies, such as the First-Come First-Serve (FCFS) policy (Gupta2011 and Gupta2010). There have also been attempts at exact solutions for multiserver queues using matrix analytic methods, but the state of the art is limited to FCFS scheduling and scheduling policies with no more than two priority classes (Sleptchenko2005).
@@ 1.2. Our contributions
In this paper, we present the first nontrivial lower bounds on the mean response time of any $M/G/k$ system.
In the prior literature, only two lower bounds on mean response time for the $M/G/k$ queue under arbitrary scheduling policies have appeared. These bounds are the mean service time and the resource-pooled single-server SRPT response time (see Section 5).
In this paper, we introduce three new lower bounds that significantly tighten these prior bounds: the MixEx bound, the ISQ bound, and the ISQ-Recycling bound, see Section 3.3 for our terminology regarding recycling. These bounds progressively improve upon one another and their degree of improvement are numerically explored in detail in Section 12. Our bounds hold for any arbitrary scheduling policy and hold across all system loads.
To derive these bounds, we introduce a novel single-server variable-speed queue, the Increasing Speed Queue (ISQ-$k$). ISQ-$k$ allows us to bridge the multiserver and single-server queues in this challenging intermediate-load environment. Moreover, we develop a novel DiffeDrift extension to the drift method to analyze variable speed queues, including ISQ-$k$.
To lower bound mean response time, we start by proving that the mean total work of the ISQ-$k$ lower bounds the mean total work of any $M/G/k$ system under arbitrary scheduling policies and at all loads (see Section 6). We build on this result to lower bound the mean relevant work of $M/G/k$ systems under arbitrary scheduling policies.
Our analysis of the ISQ-$k$ system is based on the drift method, which relies on the careful selection of test functions (Eryilmaz2012). However, most prior applications of drift methods either use standard test functions that are not customized to the system or employ rescaling techniques to eliminate variable rates (Braverman2017 and Hurtado-Lange2020). Neither of these methods can accommodate complex variable work completion rates and give a tight analysis of the ISQ-$k$ system.
@@
Figure 1. $M/G/2$ mean response time under SRPT-2, compared against naive lower bounds (M/G/1M/G/1/SRPT and $M/G/\infty$, and against our lower bounds. See Section 5 for the definition of MixEX, ISQ and ISQ-Recycling bounds.
@@ 1.3. Improvement upon prior results
We now show the massive improvement of our lower bounds over the prior naive lower bounds, all of these bounds. Consider a two-server setting with an exponential job size distribution, as shown in Figure 1. The blue line represents the mean response time of the resource pooled SRPT system, and the green line represents the mean service time bound. The empirical mean response time of $M/G/2$-SRPT is given by the dashed orange curve. While SRPT-2 is known to not be optimal in this setting (Grosof2026), it gives a useful sense of where the optimal policy may lie. Both lower bounds are observed to be loose in the intermediate load regime.
The MixEx bound given by the red curve, represents our novel combination of the mean service time bound and the resource-pooled SRPT bound via the WINE method (see Section 5). This bound significantly improves upon the naive bounds, particularly in the intermediate regime. The purple curve is the ISQ bound, which incorporates our novel ISQ-$k$ system. Finally, we call our strongest lower bound, represented by the brown curve in Figure 1, the the ISQ-Recycling bound.
To quantify the degree of improvement of our novel lower bounds, we introduce a key performance metric, the Uncertainty Improvement Ratio (UIR), which we define in Definition 12.1. UIR measures the fraction of the uncertainty region closed by a given bound relative to the best prior lower bound. More concretely, the uncertainty region is defined as the gap between the simulated SRPT-$k$ response time and the strongest available prior bound. A new bound achieves a UIR equal to the percentage of this gap that it eliminates.
With the help of ISQ-$k$, our ISQ-Recycling bound provides a substantial improvement over our MixEx, which itself represents a significant advance over the naive bounds. In Figure 2 with exponential job size distribution, MixEx achieves a maximum UIR of about 50% relative to the naive bounds. Building on this, the ISQ-Recycling bound attains a UIR exceeding 30% relative to MixEx in the same setting and for a total UIR of 62% relative to the naive bound, while maintaining meaningful improvement across the entire medium-load range ($\rho=0.4$ to $0.95$). Moreover, in the low variability setting, the ISQ-Recycling bound attains a notably high UIR of 70% relative to MixEx bound (see Section 12.3).
By contrast, the recently introduced SEK policy (Grosof2026), which was proven to outperform SRPT-$k$ at all load levels and aims to tighten upper bounds through improved simulated response times, achieves only a single-digit empirical UIR of 3% (see Appendix D). This highlights the scale of our improvement: while prior advances in upper bounds have yielded only modest gains, our framework achieves order-of-magnitude more UIR than new upper bounds, substantially narrowing the uncertainty region (see Section 12.4).
Figure 2. UIR relative to MixEx bound of the ISQ and ISQ-Recycling bound in the $M/G/2$ setting with $\text{Exp}(1)$ job sizes. See Section 5 for definitions of ISQ and ISQ-Recycling bounds.
Our theoretical framework applies to any job size distribution, any number of servers $k\geq 2$, and any load in the $M/G/k$ system, as well as any scheduling policy.
@@ 1.4. Outline of paper
  • Section 2: We discuss prior work on multiserver scheduling and the drift method.
@@
  • Section 4: We present our main results and provide a proof outline.
  • Section 5: We introduce the WINE-based framework for lower bounding the mean response time of the $M/G/k$ system and combine it with existing lower bounds.
  • Section 6: We establish results comparing the work in the ISQ-$k$ system to that in the $M/G/k$ system.
@@
  • Section 8: We start by proving our main results in the $k=2$ server case, lower bounding the $M/G/2$ under arbitrary scheduling policies before generalizing our results to the $k$-server setting.
  • Section 9: We introduce the DiffeDrift method, which we use to derive the ISQ-2 test functions, and generalize our results to the $k$ server system.
  • Section 10: We prove our main results in the general $k\geq 3$ case, lower bounding $M/G/k$ mean response time under arbitrary scheduling policies.
  • Section 11: We discuss how our framework can be extended to settings with unknown job sizes and to cases where job sizes are estimated. We also discuss computational considerations for large $k$.
@@ 2. Prior Work
2.1. Shortest remaining processing time
As discussed in the introduction, the Shortest-Remaining Processing Time (SRPT) policy is optimal in minimizing the mean response time in the $M/G/1$ queue (Schrage1968Proof and Schrage1966). In the multiserver setting, SRPT-$k$ is heavy-traffic optimal and provides an upper bound on the mean response time that is tight at very high loads (Grosof2018).
At the opposite extreme, when there are no arrivals, it is well known that the Shortest Job First (SJF) policy is the optimal nonpreemptive policy in the multiserver setting. Furthermore, McNaughton1959 proves that no preemptive policy can outperform the optimal nonpreemptive policy. Since SRPT reduces to SJF when there are no arrivals, SRPT is also optimal in the multiserver setting.
However, SRPT-$k$ is not always the optimal policy across all load conditions. Notably, the SEK policy, recently introduced by Grosof2026, empirically outperforms the response time of SRPT-$k$ under certain conditions. This raises the question of how much further improvement beyond SRPT-$k$ is possible.
The focus of this paper is on establishing lower bounds for the mean response time of $M/G/k$ systems under arbitrary policies, which has been noted as an open problem (Grosof2019Open).
@@ 2.2. Multiserver mean response time analysis
The analysis of multiserver FCFS systems, particularly in the context of $M/G/k$-FCFS queues, is relatively well understood. Early work by Kingman1970 and Daley1998 provides bounds on the mean response time based on the first two moments of the job size distribution. Harchol-Balter2005 conduct an exact analysis of the mean response time in $M/Ph/k$-FCFS systems. Since any arbitrary distribution can be closely approximated by a phase-type distribution, their results offer a useful approximation for mean response time in $M/G/k$-FCFS systems.
More recently, Gupta2011 demonstrate how utilizing the full sequence of moments of the job size distribution can yield bounds on mean response time. Li2024 extend Kingman’s bounds (Kingman1962) to the multiserver $GI/GI/k$ queue and provide the first simple and explicit bounds for mean waiting time that scale with $1/(1-\rho)$. These results allow for bounds on the mean response time in $M/G/k$-FCFS queues across all load levels. In contrast, our work considers a more general setting by allowing for any arbitrary scheduling policies and focuses on establishing lower bounds.
2.3. Relevant work analysis
Relevant work is the amount of work in the system that is prioritised ahead of a given job. Relevant-work analysis has been widely used to study the response time of many scheduling policies, such as Preemptive Shortest Job First (PSJF) and Least Attained Service (LAS) (Harchol-Balter2013).
Relevant work analysis can be applied to both single server and multiserver systems. The Work Integral Number Equality (WINE) formula, originally introduced by Scully2020 and discussed in detail in Section 5, provides a new method for computing mean response time using SRPT-based relevant work. This formula holds for any queueing system under any scheduling policy not just the SRPT-based policy. Our approach builds on this framework, using WINE as the foundation for developing lower bounds on the mean response time.
Relevant work analysis was specifically developed for index policies. These are policies that assign each job a numerical index computed from the job’s own state, and select the job with the best index. Scully2018 introduce the terms “recycling” and “recycled jobs” to describe jobs whose original index exceeds a cutoff and later decreases below that cutoff, in order to keep track of more complex index functions such as the Gittins index. Since then, this concept and terminology have been used in several subsequent works, including the Gittins-$k$ policy (Scully2020) and ServerFilling-SRPT (Grosof2023Optimal). In our setting, we preserve this terminology and refer to jobs with an initial size larger than $x$ that have received service to bring their remaining size below $x$ as recycled jobs and such events as recyclings.
Finally, Scully2020 prove that the Gittins-$k$ policy is nearly optimal in the $M/G/k$ under extremely general conditions. The Gittins-$k$ policy coincides with SRPT-$k$ when there is perfect information about job sizes, which is the setting we study in this paper.